# Positive microlocal holonomies are globally regular

### v1

@@ 4.2. Explicit description of $\operatorname{HH}_{0}$ near a relative cycle
Let $\Lambda\subset(T^{*}_{\infty}\mathbb{R}^{2},\xi_{\text{st}})$ be a Legendrian link with basepoints $T$, and $L\subset(T^{*}\mathbb{R}^{2},\lambda_{\text{st}})$ a Lagrangian filling. Let $\mathscr{D}=\{\gamma_{1},\ldots,\gamma_{b_{1}(L)}\}$ be an $\mathbb{L}$-compressing system and $\eta:[0,1]\longrightarrow(L,\Lambda\setminus T)$ a relative cycle. Let $\mathbb{L}:=L\cup\mathscr{D}^{*}$, where $\mathscr{D}^{*}=D_{1}\cup\ldots\cup D_{b_{1}(L)}$ are the Lagrangian disks $D_{i}$ associated to $\mathscr{D}$, each $D_{i}$ with boundary $\partial D_{i}=\gamma_{i}$. There exists an open neighborhood ${\mathcal{O}}{\it p}(\eta)\subset\mathbb{L}$ which is of the form $(0,1)\times I^{\sigma}_{p_{1},\ldots,p_{k}}$ where
$I^{\sigma}_{p_{1},\ldots,p_{k}}:=I\cup I_{p_{1}}^{\sigma_{1}}\cup\ldots I_{p_{k}}^{\sigma_{k}}\subset[0,1]\times[-1,1],$
@@
where $\sigma=(\sigma_{1},\ldots,\sigma_{k})\in\{\pm\}^{k}$ is a choice of signs and $p_{i}\in(0,1)$ for $i\in[1,k]$.
See Figure 2 (left) for an example. Note that $I^{\sigma}_{p_{1},\ldots,p_{k}}$ can be understood as a 1-dimensional (arboreal) Lagrangian skeleton. The symplectic structure can be taken to be that of the cotangent bundle $(T^{*}I,\lambda_{st})$ modified accordingly so that the (positive or negative) conormals of the points $p_{i}\in I$ are part of the Lagrangian skeleton, cf. [Sta18, Section 3]. For a positive relative cycles we have all the signs of $\sigma$ being positive, as in Figure 2 (left). In other words, for a positive relative cycle, a neighborhood ${\mathcal{O}}{\it p}(\eta)\subset\mathbb{L}$ of $\eta\subset\mathbb{L}$ retracts to $\mathbb{K}_{n}([0,1])$ – as in Definition 3.7 – for some $n$.
@@
By [Nad17, Theorem 1.8], $\operatorname{Sh}^{c}_{\mathbb{K}_{n}}(S^{1})$ is equivalent to the dg-category of representations of the path algebra $k\langle Q_{K}\rangle$ of the cyclic quiver $Q_{K}$ with vertices $\{v_{1},\ldots,v_{n}\}$. Confer Example 3.8 and see Figure 2 (right) for the case of the 5-cycle quiver, $n=5$. For each point in $K$, this path algebra contains a unique arrow $a_{i}$ from $v_{i}$ to $v_{i+1}$. (Cyclically understood, i.e. there is an arrow from $v_{n}$ to $v_{1}$.) In Figure 2, vertices are depicted in green and arrows in orange. Both such vertices $v_{i},v_{i+1}$, come equipped with idempotents $e_{i},e_{i+1}$ that act as identities at the sources and targets of arrows, each corresponding to the constant path at the corresponding vertex, see e.g. [DW05]. The path algebra $k\langle Q_{K}\rangle$ is linearly generated by these elements $a_{i},e_{i}$, $i\in[1,n]$, and the algebra structure is given by concatenation of paths. In particular, we have $[e_{i},a_{i}]=a_{i}e_{i}-e_{i}a_{i}=a_{i}e_{i}=a_{i}$ for the commutators if $\kappa\geq 2$. Geometrically, each vertex corresponds to the microlocal stalk of a sheaf at each of the intervals in $S^{1}\setminus K$ and the arrows capture the microlocal parallel transport between them.
@@
For any algebra $R$, there is a vector space isomorphism $\operatorname{HH}_{0}(R,R)\cong R/[R,R]$, where the trace space $R/[R,R]$ is the vector space obtained by quotienting $R$ (as a vector space) by the vector space generated by all the commutators in $R$. For the cyclic quiver $Q_{K}$, we claim that $\operatorname{HH}_{0}(k\langle Q_{K}\rangle)\cong k[\rho]\oplus k^{n}$, where $\rho\in k\langle Q_{K}\rangle$ is the cycle element $a_{1}a_{2}\ldots a_{n}$ and the summand $k^{n}$ is spanned by the idempotents. Indeed, any element $\gamma\in k\langle Q_{K}\rangle$ corresponding to a path in $Q_{K}$ whose source vertex $s(\gamma)$ is different from its target vertex $t(\gamma)$ is contained in $[k\langle Q_{K}\rangle,k\langle Q_{K}\rangle]$ as $[e_{s(\gamma)},\gamma]=\gamma$. Similarly, the cycle $a_{1}a_{2}\ldots a_{n}$ is equivalent to any cyclic permutations $a_{i}a_{i+1}\ldots a_{i-1}$. Therefore $\operatorname{HH}_{0}(k\langle Q_{K}\rangle)\cong k[\rho]\oplus k\langle e_{1},\ldots,e_{n}\rangle$, as claimed.
[…]
@@
${\displaystyle\prod\nolimits_{i=1}^{2}\operatorname{\mbox{Mod}}(k)}$${\operatorname{Sh}_{\mathbb{K}}(D^{1})}$${\operatorname{\mbox{Mod}}(k)}$$\scriptstyle{(i_{\partial D^{1}}^{*})^{\ell}}$$\scriptstyle{\Delta^{\ell}}$
@@
Then there is an isomorphism $\operatorname{HH}_{0}(\operatorname{Sh}^{c}_{\mathbb{K}_{n},\partial D^{1}}(D^{1}))\cong k[\rho]\oplus k^{n}$, where $\rho\in{H}_{1}(D^{1},\partial D^{1},\mathbb{Z})$ is the positive generator corresponding to the relative 1-cycle given by $(D^{1},\partial D^{1})$ under under $\operatorname{HH}_{0}(\operatorname{Sh}^{c}_{\mathbb{K}_{n},\partial D^{1}}(D^{1}))\longrightarrow\operatorname{HH}_{0}(\mathrm{Loc}^{c}(D^{1},\partial D^{1}))$.
@@
Note that there is an isomorphism $\operatorname{Sh}_{\mathbb{K}_{n}(S^{1})}(S^{1})\cong\operatorname{Sh}_{\mathbb{K}_{n}(D^{1}),\partial D^{1}}(D^{1})$ as follows. Let $N\subset S^{1}\setminus K$ be a closed interval. Then we have a homotopy colimit diagram
@@ 4.3. Computation of $\operatorname{HO}$ map near a relative cycle
classifies a functor $\varphi_{x}:\operatorname{Perf}(R)\longrightarrow\operatorname{Perf}(k)$ corepresented by $x^{!}:=R^{!}\otimes_{R}\varphi_{x}^{r}(k)$, i.e. $\varphi_{x}\simeq\mathrm{Hom}_{R}(x^{!},-)$. Let $(V,\rho)$, $\rho:R\longrightarrow\mathrm{End}_{k}(V)$, be the underlying $R$-module of $\varphi_{x}^{r}(k)$ and write $x:=\varphi_{x}^{r}(k)$ to ease notation, as in [BD21, Example 3.7]. Note that $V$ is a perfect $k$-module and thus a finite-dimensional $k$-vector space. Hochschild 0-chains are given by $\operatorname{HH}_{0}(\operatorname{Perf}(R))\cong\operatorname{HH}_{0}(R)\cong R/[R,R]$, the trace space of $R$, cf. [McC94, Prop. 2.4.3]. That is, $\operatorname{HH}_{0}(\operatorname{Perf}(R))$ is the vector space spanned by the elements of $R$ quotiented by the subvector space spanned by commutators. Given $r\in R$, we let $[r]\in R/[R,R]$ denote its associated Hochschild 0-chain. The following result computes $\operatorname{HO}([r])$:
@@ 4.4. Corestriction maps and Hochschild homology
[…]
@@
[…]
@@
    Map the cycle $h_{\eta}\in\operatorname{HH}_{0}(\operatorname{\operatorname{\mu Sh}}_{\mathbb{K},\mathfrak{t}}(\mathbb{K},\mathfrak{t}))$ to $\operatorname{HH}_{0}(\operatorname{\operatorname{\mu Sh}}_{\mathbb{L},\mathfrak{t}}(\mathbb{L},\mathfrak{t}))$ via the corestriction functor $\rho_{!}$ induced by the immersion $\mathbb{K}\looparrowright\mathbb{L}$. This uses Corollary 4.10 in Subsection 4.4.
For Step (1), the cycle $h_{\eta}\in\operatorname{HH}_{0}(\operatorname{\operatorname{\mu Sh}}_{\mathbb{K},\mathfrak{t}}(\mathbb{K},\mathfrak{t}))$ is defined as follows. By a pointed version of Example 3.8, cf. Prop. 3.18, $\operatorname{\operatorname{\mu Sh}}_{\mathbb{K},\mathfrak{t}}(\mathbb{K},\mathfrak{t})$ can be identified with $\operatorname{Sh}^{c}_{\mathbb{K},\mathfrak{t}}(\eta)\cong\operatorname{Sh}^{c}_{\mathbb{K},\mathfrak{t}}(D^{1})$, where $\mathbb{K}\subset T^{*}\eta$ is considered as a relative Lagrangian skeleton and we are identifying $\eta\cong D^{1}$. Since $\eta\cap\mathfrak{t}=\partial\eta$, Corollary 4.5 implies

### v2

@@ 4.2. Explicit description of $\operatorname{HH}_{0}$ near a relative cycle
Let $\Lambda\subset(T^{*}_{\infty}\mathbb{R}^{2},\xi_{\text{st}})$ be a Legendrian link with basepoints $T$, and $L\subset(T^{*}\mathbb{R}^{2},\lambda_{\text{st}})$ a Lagrangian filling. Let $\mathscr{D}=\{\gamma_{1},\ldots,\gamma_{b_{1}(L)}\}$ be an $\mathbb{L}$-compressing system and $\eta:[0,1]\longrightarrow(L,\Lambda\setminus T)$ a relative cycle. Let $\mathbb{L}:=L\cup\mathscr{D}^{*}$, where $\mathscr{D}^{*}=D_{1}\cup\ldots\cup D_{b_{1}(L)}$ are the Lagrangian disks $D_{i}$ associated to $\mathscr{D}$, each $D_{i}$ with boundary $\partial D_{i}=\gamma_{i}$. There exists an open neighborhood ${\mathcal{O}}{\it p}(\eta)\subset\mathbb{L}$ which is of the form $(0,1)\times I^{\sigma}_{p_{1},\ldots,p_{n}}$ where
$I^{\sigma}_{p_{1},\ldots,p_{n}}:=I\cup I_{p_{1}}^{\sigma_{1}}\cup\ldots I_{p_{n}}^{\sigma_{n}}\subset[0,1]\times[-1,1],$
@@
where $\sigma=(\sigma_{1},\ldots,\sigma_{n})\in\{\pm\}^{n}$ is a choice of signs and $p_{i}\in(0,1)$ for $i\in[1,n]$.
See Figure 2 (left) for an example. Note that $I^{\sigma}_{p_{1},\ldots,p_{n}}$ can be understood as a 1-dimensional (arboreal) Lagrangian skeleton. The symplectic structure can be taken to be that of the cotangent bundle $(T^{*}I,\lambda_{st})$ modified accordingly so that the (positive or negative) conormals of the points $p_{i}\in I$ are part of the Lagrangian skeleton, cf. [STA18, Section 3]. For a positive relative cycles we have all the signs of $\sigma$ being positive, as in Figure 2 (left). In other words, for a positive relative cycle, a neighborhood ${\mathcal{O}}{\it p}(\eta)\subset\mathbb{L}$ of $\eta\subset\mathbb{L}$ retracts to $\mathbb{K}_{n}([0,1])$ – as in Definition 3.9 – for some $n$.
@@
By [NAD17, Theorem 1.8], $\operatorname{Sh}^{c}_{\mathbb{K}_{n}}(S^{1})$ is equivalent to the dg-category of representations of the path algebra $k\langle Q_{K}\rangle$ of the cyclic quiver $Q_{K}$ with vertices $\{v_{1},\ldots,v_{n}\}$. Confer Example 3.10 and see Figure 2 (right) for the case of the 5-cycle quiver, $n=5$. For each point in $K$, this path algebra contains a unique arrow $a_{i}$ from $v_{i}$ to $v_{i+1}$. (Cyclically understood, i.e. there is an arrow from $v_{n}$ to $v_{1}$.) In Figure 2, vertices are depicted in green and arrows in orange. Both such vertices $v_{i},v_{i+1}$, come equipped with idempotents $e_{i},e_{i+1}$ that act as identities at the sources and targets of arrows, each corresponding to the constant path at the corresponding vertex, see e.g. [DW05]. The path algebra $k\langle Q_{K}\rangle$ is linearly generated by these elements $a_{i},e_{i}$, $i\in[1,n]$, and the algebra structure is given by concatenation of paths. In particular, we have $e_{i},a_{i}]=a_{i}e_{i}-e_{i}a_{i}=a_{i}e_{i}=a_{i}$ for the commutators if $n\geq 2$. Geometrically, each vertex corresponds to the microlocal stalk of a sheaf at each of the intervals in $S^{1}\setminus K$ and the arrows capture the microlocal parallel transport between them.
@@
For any algebra $R$, there is a vector space isomorphism $\operatorname{HH}_{0}(R,R)\cong R/[R,R]$, where the trace space $R/[R,R]$ is the vector space obtained by quotienting $R$ (as a vector space) by the vector space generated by all the commutators in $R$. For the cyclic quiver $Q_{K}$, we claim that $\operatorname{HH}_{0}(k\langle Q_{K}\rangle)\cong k[\rho]\oplus k^{n}$, where $\rho\in k\langle Q_{K}\rangle$ is the cycle element $a_{1}a_{2}\ldots a_{n}$ and the summand $k^{n}$ is spanned by the idempotents. Indeed, any element $\gamma\in k\langle Q_{K}\rangle$ corresponding to a path in $Q_{K}$ whose source vertex $s(\gamma)$ is different from its target vertex $t(\gamma)$ is contained in $k\langle Q_{K}\rangle,k\langle Q_{K}\rangle]$ as $e_{s(\gamma)},\gamma]=\gamma$. Similarly, the cycle $a_{1}a_{2}\ldots a_{n}$ is equivalent to any cyclic permutations $a_{i}a_{i+1}\ldots a_{i-1}$. Therefore $\operatorname{HH}_{0}(k\langle Q_{K}\rangle)\cong k[\rho]\oplus k\langle e_{1},\ldots,e_{n}\rangle$, as claimed.
Finally, we note that the equivalence from [NAD17, Theorem 1.8] sends the quiver cycle $\rho$ to a homological representative $\rho$ of $S^{1}\subset T^{*}S^{1}$. In fact, consider the localization functor $\operatorname{Sh}^{c}_{\mathbb{K}_{n}}(S^{1})\longrightarrow\mathrm{Loc}^{c}(S^{1})$ given by Corollary 3.21. Let $Q_{0}$ be the quiver with one vertex and a single loop $\rho$. Then $\mathrm{Loc}^{c}(S^{1})$ is equivalent to the dg-category of representations of the path algebra localized along the loop $k\left<Q_{0}\right>_{\rho}$. Using [NAD17, Section 4.2], the localization functor $\operatorname{Sh}^{c}_{\mathbb{K}_{n}}(S^{1})\longrightarrow\mathrm{Loc}^{c}(S^{1})$ can be described by the localization morphism on path algebras $k\left<Q_{K}\right>\to k\left<Q_{0}\right>\to k\left<Q_{0}\right>_{\rho}$ which localizes $a_{1},a_{2},\dots,a_{n}$ and sends the product $a_{1}a_{2}\dots a_{n}$ to the invertible cycle $\rho$. Therefore, on Hochschild homology, we have
@@
$\prod\nolimits_{i=1}^{2}\operatorname{\mbox{Mod}}(k)$$\operatorname{Sh}_{\mathbb{K}}(D^{1})$$\operatorname{\mbox{Mod}}(k)$$\scriptstyle{(i_{\partial D^{1}}^{*})^{\ell}}$$\scriptstyle{\Delta^{\ell}}$
@@
Then there is an isomorphism $\operatorname{HH}_{0}(\operatorname{Sh}^{c}_{\mathbb{K}_{n},\partial D^{1}}(D^{1}))\cong k[\rho]\oplus k^{n}$, where $\rho\in{H}_{1}(D^{1},\partial D^{1},\mathbb{Z})$ is the positive generator corresponding to the relative 1-cycle given by $(D^{1},\partial D^{1})$ under $\operatorname{HH}_{0}(\operatorname{Sh}^{c}_{\mathbb{K}_{n},\partial D^{1}}(D^{1}))\longrightarrow\operatorname{HH}_{0}(\mathrm{Loc}^{c}(D^{1},\partial D^{1}))$.
@@
Note that there is an equivalence $\operatorname{Sh}_{\mathbb{K}_{n}(S^{1})}(S^{1})\cong\operatorname{Sh}_{\mathbb{K}_{n}(D^{1}),\partial D^{1}}(D^{1})$ as follows. Let $N\subset S^{1}\setminus K$ be a closed interval. Then we have a homotopy colimit diagram
@@ 4.3. Computation of $\operatorname{HO}$ map near a relative cycle
classifies a functor $\varphi_{x}:\operatorname{Perf}(R)\longrightarrow\operatorname{Perf}(k)$ corepresented by $x^{!}:=R^{!}\otimes_{R}\varphi_{x}^{r}(k)$, i.e. $\varphi_{x}\simeq\mathrm{Hom}_{R}(x^{!},-)$. Let $(V,\rho)$, $\rho:R\longrightarrow\mathrm{End}_{k}(V)$, be the underlying $R$-module of $\varphi_{x}^{r}(k)$ and write $x:=\varphi_{x}^{r}(k)$ to ease notation, as in [BD21, Example 3.7]. Note that $V$ is a perfect $k$-module and thus a finite-dimensional $k$-vector space. Hochschild 0-chains are given by $\operatorname{HH}_{0}(\operatorname{Perf}(R))\cong\operatorname{HH}_{0}(R)\cong R/[R,R]$, the trace space of $R$, cf. [MCC94, Prop. 2.4.3]. That is, $\operatorname{HH}_{0}(\operatorname{Perf}(R))$ is the vector space spanned by the elements of $R$ quotiented by the subvector space spanned by commutators. Given $r\in R$, we let $r]\in R/[R,R]$ denote its associated Hochschild 0-chain. The following result computes $\operatorname{HO}([r])$:
@@ 4.4. Corestriction maps and Hochschild homology
(4.8) $\operatorname{HH}_{*}(\operatorname{\operatorname{\mu Sh}}_{\mathbb{K},\mathfrak{t}}^{c}(\mathbb{K},\mathfrak{t}))$$\operatorname{HH}_{*}(\operatorname{\operatorname{\mu Sh}}_{\mathbb{L},\mathfrak{t}}^{c}(\mathbb{L},\mathfrak{t}))$$\Gamma(\mathfrak{M}(\operatorname{\operatorname{\mu Sh}}_{\mathbb{K},\mathfrak{t}}^{c}(\mathbb{K},\mathfrak{t})),\mathcal{O}_{\mathfrak{M}(\operatorname{\operatorname{\mu Sh}}_{\mathbb{K},\mathfrak{t}}^{c}(\mathbb{K},\mathfrak{t}))})$$\Gamma(\mathfrak{M}(\operatorname{\operatorname{\mu Sh}}_{\mathbb{L},\mathfrak{t}}^{c}(\mathbb{L},\mathfrak{t})),\mathcal{O}_{\mathfrak{M}(\operatorname{\operatorname{\mu Sh}}_{\mathbb{L},\mathfrak{t}}^{c}(\mathbb{L},\mathfrak{t}))})$$\scriptstyle{\operatorname{HH}_{*}(\rho_{!})}$$\scriptstyle{\operatorname{HO}_{\mathbb{K}}}$$\scriptstyle{\operatorname{HO}_{\mathbb{L}}}$
@@
(4.9) $\operatorname{HH}_{*}(\mathrm{Loc}^{c}(\eta,\mathfrak{t}\cap\eta))$$\operatorname{HH}_{*}(\mathrm{Loc}^{c}(L,\mathfrak{t}))$$\Gamma(\mathfrak{M}(\mathrm{Loc}^{c}(\eta,\mathfrak{t}\cap\eta)),\mathcal{O}_{\mathfrak{M}(\mathrm{Loc}^{c}(\eta,\mathfrak{t}\cap\eta))})$$\Gamma(\mathfrak{M}(\mathrm{Loc}^{c}(L,\mathfrak{t})),\mathcal{O}_{\mathfrak{M}(\mathrm{Loc}^{c}(L,\mathfrak{t}))}).$$\scriptstyle{\operatorname{HO}_{\eta}}$$\scriptstyle{\operatorname{HO}_{L}}$
@@
    Map the cycle $h_{\eta}\in\operatorname{HH}_{0}(\operatorname{\operatorname{\mu Sh}}_{\mathbb{K},\mathfrak{t}}(\mathbb{K},\mathfrak{t}))$ to $\operatorname{HH}_{0}(\operatorname{\operatorname{\mu Sh}}_{\mathbb{L},\mathfrak{t}}(\mathbb{L},\mathfrak{t}))$ via the corestriction functor $\rho_{!}$ induced by the immersion $\mathbb{K}\looparrowright\mathbb{L}$. This uses Corollary 4.9 in Subsection $4.$
For Step (1), the cycle $h_{\eta}\in\operatorname{HH}_{0}(\operatorname{\operatorname{\mu Sh}}_{\mathbb{K},\mathfrak{t}}(\mathbb{K},\mathfrak{t}))$ is defined as follows. By a pointed version of Example 3.10, $\operatorname{\operatorname{\mu Sh}}_{\mathbb{K},\mathfrak{t}}(\mathbb{K},\mathfrak{t})$ can be identified with $\operatorname{Sh}^{c}_{\mathbb{K},\mathfrak{t}}(\eta)\cong\operatorname{Sh}^{c}_{\mathbb{K},\mathfrak{t}}(D^{1})$, where $\mathbb{K}\subset T^{*}\eta$ is considered as a relative Lagrangian skeleton and we are identifying $\eta\cong D^{1}$. Since $\eta\cap\mathfrak{t}=\partial\eta$, Corollary 4.4 implies
@@
4.6. A comment on generality
We have stated and proven 1.1 and 1.2 for Legendrian links $\Lambda\subset(T^{*}_{\infty}\mathbb{R}^{2},\xi_{\text{st}})$. This specific framework is the most studied in the literature, see e.g. [CG22, CG24, CGG+25, CW24, CZ22, EHK16]. By the Darboux theorem, this case needs to be understood if one aims at studying Lagrangian fillings of an arbitrary Legendrian link $\Lambda\subset(\partial W,\ker\lambda)$ in the contact boundary of a Weinstein 4-manifold $(W,\lambda)$. Following a suggestion of the referee, we briefly sketch a generalization of 1.1, as follows.
The proof of 1.1, as written, works almost identically if one considers embedded exact Lagrangian fillings $L\subset(W,\lambda_{\text{st}})$ of a Legendrian link $\Lambda\subset(\partial W,\ker\lambda_{\text{st}})$. The main modifications are:
  1. (1)
    The definition of an $\mathbb{L}$-compressing system $\mathscr{D}$ for $L$ generalizes as follows: an $\mathbb{L}$-compressing system $\mathscr{D}$ for $L\subset(W,\lambda_{\text{st}})$ is still a collection of Lagrangian disks in $W$ with boundary on $L$, properly embedded on its complement $W\setminus L$, with the condition that the arboreal Lagrangian given by the union $\mathbb{L}:=L\cup_{\partial\mathscr{D}}\mathscr{D}$ must now be a relative Lagrangian skeleton for the Weinstein pair given by the Weinstein 4-manifold $(W,\lambda_{\text{st}})$ and the Weinstein ribbon of $\Lambda$. See e.g. [ELI18, Section 2] for details on Weinstein pairs.
    Note that the Kashiwara-Schapira stack $\operatorname{\operatorname{\mu Sh}}_{\mathbb{L}}$ from Section 3 is still defined, cf.  e.g. [NAD16, Section 3.4], and the results on $\operatorname{\operatorname{\mu Sh}}_{\mathbb{L}}$ and corestriction functors remain the same. Similarly, the pointed versions $\operatorname{\operatorname{\mu Sh}}_{\mathbb{L},\mathfrak{t}}(\mathbb{L})$ and $\mathfrak{M}(\mathbb{L},\mathfrak{t})$ of the category $\operatorname{\operatorname{\mu Sh}}_{\mathbb{L}}(\mathbb{L})$ and its moduli stack $\mathfrak{M}(\mathbb{L})$ are defined via the same homotopy colimits and limits.
  2. (2)
    The condition of the relative 1-cycle $\eta$ being $\mathscr{D}$-positive can be generalized to the condition that a small enough neighborhood of $\eta\subset L$ in the Lagrangian skeleton $\mathbb{L}=L\cup_{\partial\mathscr{D}}\mathscr{D}$ retracts to the model in Section 4.2, i.e. to a horizonal line (or circle) with a sequence of half-rays attached to it all pointing in the upwards direction, as in Fig. 2 (left). We still refer to such a relative 1-cycle $\eta$ as a $\mathscr{D}$-positive cycle.
The same ingredients as in Section 1.4, with the modified concepts above, leads to a generalized version of 1.1. The statement reads as follows:
Theorem 4.11 (Generalized extension result).
Let $\Lambda\subset(\partial W,\ker\lambda_{\text{st}})$ be a $(\mathfrak{t},T)$-pointed Legendrian link in the contact boundary of a Weinstein 4-manifold $(W,\lambda_{\text{st}})$ and $L\subset(W,\lambda_{\text{st}})$ an embedded exact Lagrangian filling of $\Lambda$ endowed with an $\mathbb{L}$-compressing system $\mathscr{D}$.
Then, for any $\mathscr{D}$-positive relative 1-cycle $\eta$ in the relative pair $(L,T)$, there exists a Hochschild 0-cycle $H_{\eta}\in\operatorname{HH}_{0}(\operatorname{\operatorname{\mu Sh}}_{\mathbb{L},\mathfrak{t}}(\mathbb{L}))$ whose associated regular function $\operatorname{HO}(H_{\eta})\in\mathrm{H}^{0}\Gamma(\mathfrak{M}({\mathbb{L},\mathfrak{t}}),\mathcal{O}_{\mathfrak{M}({\mathbb{L},\mathfrak{t}})})$ coincides with the trace of the microlocal merodromy along $\eta$ when restricted to ${T}(L)\subset\mathfrak{M}({\mathbb{L},\mathfrak{t}})$.∎
Note that the symplectic invariance and properties of the category $\operatorname{\operatorname{\mu Sh}}_{\mathbb{L}}(\mathbb{L})$ for an arbitrary Lagrangian skeleton $\mathbb{L}$ are still being explored, see e.g. [NS20]. For the case of $(W,\lambda)=(T^{*}S,\lambda_{\text{st}})$ of the standard cotangent bundle of a smooth surface $S$, where $S$ is a smooth Lagrangian skeleton, the results from [GKS12] apply, see also [GUI23, Section 10], and the resulting categories are classically known to be invariant under contact isotopies and compactly supported Weinstein homotopies.